# Distributed ADMM Approach for the Power Distribution Network Reconfiguration

### v1

@@ IV-B2 Resilience
$\delta=0.1$ 140.1 132.28-143.11 1430 $/10$ 10
$\delta=1$ 156.43 143.76-163.18 1190 $/10$ 10
$\delta=10$ 172.3 162.9-180.2 871 $/10$ 10
$\delta=100$ 188.4 176.08-198.55 517 $/10$ 10
$\delta=10^{8}$ 201.76 190.65-212.62 497 $/10$ 10
(a) The evolution of $e_{k}$ in function of $k$ without fault.
(b) The evolution of $e_{k}$ in function of $k$ in the case of fault occurred in the $1200$th iteration.
Figure 4: The evolution of $e_{k}$ as a function of $k$ in scenarios with and without a fault.
@@ IV-C Centralized vs distributed
In this section, we compare Algorithm 1 with its centralized counterpart, which involves the following iterations:
@@
TABLE II: Results for centralized algorithm on the 33-bus system with different values of $\delta$.
Average Losses (kW) Min-Max Losses (kW) Average No. of Iterations Average Time (s) Feasibility Ratio
$\delta=1$ 137.8 132.09-140.87 146 8.27 $/10$ 10
$\delta=10$ 152.3 146.7-158.4 31 1.12 $/10$ 10
$\delta=100$ 166.8 152.12-188.7 14 0.32 $/10$ 10
$\delta=10^{8}$ 198.2 175.2-224 4 0.09 $/10$ 10
@@ IV-C1 Algorithm 1 Compared to [17] and [18]
where $\boldsymbol{w}^{i}\in[0,1]^{|\mathcal{A}|}$ is the relaxation of $\boldsymbol{b}^{i}$ and $\boldsymbol{u}^{i}$ is the Lagrange multiplier for the constraints $\boldsymbol{w}^{i}=\boldsymbol{b}^{i}$, $i\in\mathcal{V}$. The improvement of [18] over [17] lies in the level of the projection operator where the Douglas-Rachford splitting method was used. However, in both papers, problem (45) is treated as a projection problem of $\boldsymbol{w}^{i}_{k+1}+\boldsymbol{u}_{k}^{i}$ onto the set $\{0,1\}^{|\mathcal{A}|}$, which does not ensure the tree structure of the solution. Instead, in the same spirit of Proposition III.1, it can be shown that problem (45) can be solved by solving an MWRAP with weights:
@@
IV-D Real case study
We conducted experiments on a real-world power grid managed by SRD, which supplied the necessary data. This grid comprises 810 buses, 8 substations, and 1 generator, as illustrated in Figure 5. The total active and reactive loads are $5.1267$ MW and $0.51$ MVAr respectively with voltage magnitude $V_{0}=20$ kV.
@@ IV-D Real case study
Figure 5: The SRD power grid. The big red node in the center of the network represents the generator, the green nodes denote the substations. The red lines are the open switches in the optimal case.
Figure 6: The evolution of $e_{k}$ as a function of $k$ for $\delta=10^{8}$ and $\boldsymbol{b}_{0}=\boldsymbol{0}_{|\mathcal{A}|}$.
The results of these experiments are detailed in Table IV. The optimality gap is calculated as:
Gap=(ADMM lossesSRD $\text{Gap}=\left(\frac{\text{ADMM losses}}{\text{SRD losses}}-1\right)\times 100\%,$
which provides a measure of the relative difference between the losses computed by Algorithm 1 and the optimal losses established by SRD. The ”ADMM losses” represent the values obtained from Algorithm 1, and the ”SRD losses” represent the optimal solution values from SRD. It’s important to note that due to the simplifications in our model, the losses derived serve as approximations rather than exact values.
TABLE IV: Results of execution of Algorithm 1 on the SRD network with different values of $\delta$.
$\delta$ Avr. Gap Min-Max. Gap (kw) Avg. No. of Iter
$1$ $24.316\%$ $15.29\%-34.67\%$ $19551$
$10$ $34.85\%$ $17.43\%-39.86\%$ $15298$
$100$ $39.4\%$ $27.12\%-54.76\%$ $12737$
$10^{8}$ $54.02\%$ $43.61\%-70.43\%$ $9874$
Table IV presents results for different values of the parameter $\delta$, illustrating the average optimality gap, the range of gap values, and the average iterations required for convergence. The average is computed for Algorithm 1 across $10$ restarts, each utilizing a distinct initial state $\boldsymbol{b}_{0}$.”
An increase in $\delta$ corresponds to a notable reduction in iterations, suggesting faster convergence. However, this acceleration comes at the cost of the solution quality, measured by the optimality gap, which diminishes as a trade-off.
We further test the resilience of our algorithm in response to real-time faults in the network. For $\delta=10^{8}$ and $\boldsymbol{b}_{0}=\boldsymbol{0}_{|\mathcal{A}|}$, Algorithm 1 took 8698 iteration to converge. A fault is introduced at iteration 8500 by deactivating a line. Figure 6 illustrates the evolution of $e_{k}$, demonstrating that our algorithm successfully adapts to the fault and continues the convergence process. In this scenario, convergence required $2\times 10^{4}$ iterations.

### v2

@@ IV-B3 Resilience
$\delta=0.1$ 140.1 132.28-143.11 1430 $10/10$
$\delta=1$ 156.43 143.76-163.18 1190 $10/10$
$\delta=10$ 172.3 162.9-180.2 871 $10/10$
$\delta=100$ 188.4 176.08-198.55 517 $10/10$
$\delta=10^{8}$ 201.76 190.65-212.62 497 $10/10$
(a) Error evolution $e_{k}$ without fault.
(b) Zoom-in on the $\ell^{2}$-norm of the 33 agents, $\|\boldsymbol{X}_{k}^{i}\|$ without fault.
(c) Error evolution $e_{k}$ with fault at iteration 1200.
(d) $\ell^{2}$-norm of the 33 agents $\|\boldsymbol{X}_{k}^{i}\|$ with fault at iteration 1200.
Figure 4: Evolution of the error $e_{k}$ and the agent norm $\|\boldsymbol{X}_{k}^{(i)}\|$ in scenarios without fault (top) and with a fault introduced at iteration 1200 (bottom). All the agents converge toward the same limit (consensus).
@@ IV-C Centralized vs distributed
In this section, we compare Algorithm 1 with its centralized counterpart proposed in [13], which involves the following iterations:
@@
TABLE II: Results for the centralized algorithm on the 33-bus system with different values of $\delta$.
$\delta=1$ 137.8 132.09–140.87 146 $10/10$
$\delta=10$ 152.3 146.7–158.4 31 $10/10$
$\delta=100$ 166.8 152.12–188.7 14 $10/10$
$\delta=10^{8}$ 198.2 175.2–224 4 $10/10$
@@ IV-C1 Algorithm 1 Compared to [19] and [20]
where $\boldsymbol{w}^{i}\in[0,1]^{|\mathcal{A}|}$ is the relaxation of $\boldsymbol{b}^{i}$ and $\boldsymbol{u}^{i}$ is the Lagrange multiplier for the constraints $\boldsymbol{w}^{i}=\boldsymbol{b}^{i}$, $i\in\mathcal{V}$. The improvement of [20] over [19] lies in the level of the projection operator where the Douglas-Rachford splitting method was used. However, in both papers, problem (48) is treated as a projection problem of $\boldsymbol{w}^{i}_{k+1}+\boldsymbol{u}_{k}^{i}$ onto the set $\{0,1\}^{|\mathcal{A}|}$, which does not ensure the tree structure of the solution. Instead, in the same spirit of Proposition III.1, it can be shown that problem (48) can be solved by solving an MWRAP with weights (see [13, Proposition 3.2]):
@@
IV-C2 Conclusion
The numerical results demonstrate that Algorithm 1 effectively achieves consensus and optimal configurations in both standard and fault scenarios. While convergence to the global optimum depends on initialization, the algorithm consistently produces feasible and near-optimal solutions. Its resilience to faults and adaptability to changing conditions confirm its robustness.
Although the distributed algorithm requires more iterations than its centralized counterpart, it offers superior scalability and supports fully decentralized execution. Furthermore, it consistently outperforms relaxation-based distributed methods [19, 20] in both feasibility and convergence reliability, owing to its exact projection step—formulated as a MWRAP which guarantees the radial solutions.
IV-D Real Case Study: SRD Network
IV-D1 Objective
The objective of this subsection is to evaluate the scalability and real-world applicability of Algorithm 1 on a large-scale operational distribution network. Specifically, we aim to assess both the convergence performance and the algorithm’s robustness when applied to realistic grid topologies and load profiles provided by an actual utility.
IV-D2 Experiment
We test Algorithm 1 on a real distribution system operated by SRD, comprising $810$ buses, $8$ substations, and $1$ generator. The total load is $5.1267$ MW and $0.51$ MVAr, with a nominal voltage magnitude of $V_{0}=20$ kV. The network topology is illustrated in Figure 5. To evaluate convergence behavior, we vary the parameter $\delta$ and compute the average optimality gap across $10$ different random initializations $\boldsymbol{b}_{0}$. The gap is measured as:
$\text{Gap}=\left(\frac{\text{ADMM losses}}{\text{SRD losses}}-1\right)\times 100\%,$
where the SRD losses represent reference values from the utility’s solution.
IV-D3 Results
Table IV summarizes the average gap, min-max gap, and average iteration count as a function of $\delta$. We observe a clear trade-off: increasing $\delta$ accelerates convergence but leads to a degradation in solution quality, with the optimality gap rising from $24.3\%$ to $54\%$ as $\delta$ increases from $1$ to $10^{8}$. Despite the model simplifications, the results demonstrate consistent convergence behavior across runs.
@@ IV-D3 Results
Figure 5: The SRD power grid. The big red node in the center represents the generator; green nodes indicate substations. Red lines show open switches in the optimal configuration.
TABLE IV: Performance of Algorithm 1 on the SRD network for various $\delta$ values. Averages computed over 10 initializations.
$\delta$ Avg. Gap (%) Min-Max Gap (%) Avg. Iterations
11 $24.32$ 15.2915.2934.6734.67 $19551$
$10$ $34.85$ 17.4317.4339.8639.86 $15298$
$100$ $39.40$ 27.1227.1254.7654.76 $12737$
$10^{8}$ $54.02$ 43.6143.6170.4370.43 $9874$
IV-D4 Resilience Evaluation
To assess real-time adaptability, a fault was injected at iteration $8500$ by deactivating a line, under the setting $\delta=10^{8}$ and $\boldsymbol{b}_{0}=\boldsymbol{0}$. As illustrated in Figure 6, the algorithm responded to this disturbance by adapting the solution trajectory and converging in approximately $2\times 10^{4}$ iterations. This showcases its capacity to reconfigure dynamically in large-scale settings.
Figure 6: Evolution of $e_{k}$ with fault injection at iteration 8500, for $\delta=10^{8}$.
IV-D5 Conclusion
These experiments confirm the scalability, resilience, and practicality of Algorithm 1 on real-world distribution networks. Despite inherent simplifications, the method achieves consistent feasibility and convergence, while adapting to dynamic changes without centralized control.